OpenStudy (anonymous):

cos(x+45)+cos(x-45)=sqrt2

OpenStudy (anonymous):

use the sum and different identities. cos(alpha+beta) and cos(alpha-beta)

OpenStudy (anonymous):

(cosxcos45-sinxsin45)+(cosxcos45+sinxsin45)=sqrt2 2cosxcos45=sqrt2 cosxcos45=sqrt2/2 (sqrt2/2)cosx=sqrt2/2 cosx=1 x=arccos1=0